OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (yrelhan4):

This is an AP with common difference -5 and first term 1.. you know the formula for sum of n terms of an AP?

OpenStudy (anonymous):

\[S{n}=\frac{ n }{ 2 }(a _{1}+a_{n})\]

OpenStudy (anonymous):

right?

OpenStudy (yrelhan4):

yeah. if you want to use this.. you need to find the twelfth term.. you know the formula for that?

OpenStudy (anonymous):

im not sure

OpenStudy (yrelhan4):

formula to find the nth term..

OpenStudy (yrelhan4):

an=a+(n-1)d

OpenStudy (anonymous):

so the formula for the nth term of an arithmetic sequence?

OpenStudy (yrelhan4):

where a is the first term, n is the number of terms.. put n=12 and find the twelfth term..

OpenStudy (yrelhan4):

yeah. a_12 = 1 + 11(-5) = 1 - 55 = -54.. so your 12th term is -54..

OpenStudy (anonymous):

where did you get -5?

OpenStudy (anonymous):

oh wait nevermind i remember

OpenStudy (anonymous):

We remark that : \[-4-1=-5\\ -9-(-4)=-5 \\-14-(-9)=-5\] The difference between every consecutive terms is -5, so this is an arithmetic progression with first term u_1=1 and the common difference d=-5. So : \[S_{12}=u_1+\cdots u_{12}=\frac{12}{2}(2u_1+11d)\]

OpenStudy (yrelhan4):

now you have a_n, a.. and you know n which is 12.. plug all of that in that sum formula.. thats the common differnce..

OpenStudy (yrelhan4):

So S_n = (12/2)*(1 + (-54) = 6*(-53) = -318.. thats your answer.

OpenStudy (anonymous):

okay, im not sure i understand fully, but thank you.

OpenStudy (yrelhan4):

Reply back if you have any problems. You're welcome.